OpenStudy (anonymous):

A tortoise is walking in the desert. It walks at a speed of 5 meters per minute for 12.5 meters. For how many minutes does it walk?

OpenStudy (anonymous):

12.5 x 5 = 62.5 minutes

OpenStudy (anonymous):

might be wrong ... possibly

OpenStudy (johnweldon1993):

So for 12.5 meters...the little guy walks 5 meters per minute.... \[\large \frac{12.5 meters}{5\frac{meters}{minute}} = 2.5minutes \] we can check it by saying...okay...if his speed was 5 meters per minute....and he walked for 2.5 minutes...does he walk 12.5 meters? \[\large 5 \times 2.5 = 12.5\] so yes it checks out :)